# Distilling circularly polarized photon emission from magnetized chiral plasmas

### v1

@@ I Introduction
High temperature and energy density are not the only extreme characteristics of the QGP produced by heavy-ion collisions. The system is, without exaggeration, rich in extremes. Since heavy ions carry electric charges, they induce extreme magnetic fields [1, 2, 3, 4, 5]. Furthermore, large values of angular momentum in non-central collisions can produce high vorticity in the resulting QGP fluid [6, 7]. Many theoretical aspects of QGP associated with strong magnetic fields and high vorticity are routinely tested in experimental studies [8, 9, 10, 11]. (For reviews, see Refs. [12, 13, 14, 15, 16].)
One aspect we would like to discuss in more detail is the effect of chiral charge density on the electromagnetic probes of QGP. A nonzero chiral charge arises from an imbalance between left-handed and right-handed quarks in the plasma. At first glance, such a state may seem too exotic as it would break local parity symmetry. However, in a hot plasma, local bubbles of chiral plasma can be produced by topological configurations of gauge fields [17]. This idea is not new, and it is taken seriously enough to motivate the challenging search for its observable signatures via the dipole and quadrupole chiral magnetic effects (CME) in multiple experiments. Unfortunately, the evidence supporting CME remains inconclusive.
@@
This paper is organized as follows. In Sec. II, we briefly present the formalism used for studying polarized photon emission from the plasma. Numerical results for the polarized photon emission are given in Sec. III. In Sec. IV, we discuss our main findings and summarize the results. Several technical derivations and auxiliary results are provided in the appendices at the end of the paper.
@@ II Formalism
To investigate the impact of a nonzero chiral charge density on photon emission from a strongly magnetized plasma, we will utilize the general framework developed in Refs. [21, 22]. Specifically, we will utilize a relationship between the emission rate and the imaginary part of the retarded photon polarization tensor [23]
@@
Assuming that the background magnetic field is sufficiently strong, the polarization tensor $\Pi_{\mu\nu}(k)$ is determined by the simplest one-loop diagram at the leading order of coupling [22]. For QED plasma, this approximation should hold true for magnetic field strengths $|eB|\gtrsim\alpha T^{2}$, where $\alpha=1/137$ denotes the fine structure constant. In the context of QGP, the requirement is $|eB|\gtrsim\alpha_{s}T^{2}$, where $\alpha_{s}$ is the strong coupling constant. Additionally, in the case of QCD, one should require that $T\gg\Lambda_{\rm QCD}$.
To separate total emission into polarized components, we use the following completeness relation:
$g^{\mu\nu}=\epsilon_{0}^{\mu}\epsilon_{0}^{\nu*}-\epsilon_{+}^{\mu}\epsilon_{+}^{\nu*}-\epsilon_{-}^{\mu}\epsilon_{-}^{\nu*}-\epsilon_{\parallel}^{\mu}\epsilon_{\parallel}^{\nu*}.$ (2)
@@
Interestingly, the extra contribution ${\cal R}^{\rm(extra)}_{\rm diff}$ to the rate, coming from $\Pi^{\rm(extra)}(k)\equiv\Pi^{(0)}(k)+\Pi^{(\parallel)}(k)$, is negative. This may seem problematic since the rates are expected to be positive definite. However, as one can verify, the total rate ${\cal R}^{\rm(extra)}_{\rm diff}$ is always positive. This implies that the extra contribution ${\cal R}^{\rm(extra)}_{\rm diff}$ should be distributed between the two physical circular polarizations, modifying the definitions of rates in Eqs. (5) and (6). For simplicity, we will use the unmodified definitions above, which should suffice to approximately extract parity-odd contributions (albeit somewhat overestimated) to the rates for circularly polarized photons.
@@
$\bar{G}_{f}(p_{0},p_{z};\mathbf{r}_{\perp})=i\frac{e^{-\mathbf{r}_{\perp}^{2}/(4\ell_{f}^{2})}}{2\pi\ell_{f}^{2}}\sum_{n=0}^{\infty}\frac{\tilde{D}_{n}^{f}(p_{0},p_{z};\mathbf{r}_{\perp})}{\mathcal{M}-2n|e_{f}B|}.$ (7)
@@
Here, $L_{n}^{\alpha}(z)$ are the generalized Laguerre polynomials [25], $\mathcal{P}_{\pm}^{f}\equiv\frac{1}{2}\left(1\pm is_{\perp}\gamma^{1}\gamma^{2}\right)$ are the spin projectors, $\mathbf{r}_{\perp}=(x,y)$ is the transverse position vector, and $\ell_{f}=1/\sqrt{|e_{f}B|}$ is the magnetic length. Here, we included a nonzero quark number and chiral chemical potentials, denoted by $\mu$ and $\mu_{5}$, respectively. By definition, $s_{\perp}=\,\mbox{sign}(e_{f}B)$ and $L_{-1}^{\alpha}(z)\equiv 0$.
@@
where $\mathcal{P}_{5}^{(\pm)}=\frac{1}{2}(1\pm\gamma^{5})$ are the chirality projectors. Note that we included both the quark-number and chiral-charge chemical potentials, $\mu$ and $\mu_{5}$, respectively. Both play important role in polarized photon emission. For the sake of simplicity, we assume that $\mu$ and $\mu_{5}$ are same for both flavors. (The generalization to the case of flavor-dependent chemical potentials is straightforward and will be briefly discussed later.)
@@
where we used the Matsubara finite-temperature formalism. By definition, the fermionic and bosonic Matsubara frequencies are $\omega_{k}=(2k+1)\pi T$ and $\Omega_{m}=2n\pi T$, respectively. Regarding other notations, $\alpha_{f}=q_{f}^{2}\alpha$ is the flavor-dependent coupling constant, and $N_{c}=3$ is the number of quark colors.
By substituting Green’s function (7) into the definition for $\Pi^{\mu\nu}(i\Omega_{m};\mathbf{k})$, calculating the Dirac traces, integrating over $r_{\perp}$ and performing the Matsubara summation, we derive the following expression for the components of the polarization tensor:
@@
where $i$ labels the photon polarizations (i.e., $i=0,+,-,\parallel$), $E_{n,p_{z},f}=\sqrt{p_{z}^{2}+2n|e_{f}B|}$ are Landau-level energies, and $n_{F}(E)=1/\left[\exp(E/T)+1\right]$ is the Fermi-Dirac distribution function. The corresponding technical details of the derivation are given in Appendices A, B, C, and D. The explicit forms of the polarization-dependent functions $\mathcal{F}^{(i)}_{s,f}(\xi,i\Omega_{m})$ are presented in Eqs. (40) – (43) in Appendix E.
@@
where $\theta(x)$ is the Heaviside step function. In Eq. (12), we also introduced two transverse momentum thresholds,
$k_{\pm}^{f}=\left|\sqrt{2n|e_{f}B|}\pm\sqrt{2n^{\prime}|e_{f}B|}\right|.$ (13)
@@
As expected, the expression in Eq. (12) closely resembles the result in Ref. [22]. It should be noted, however, that $g(n,n^{\prime})$ depends on both fermion-number and chiral chemical potentials. The most important difference appears in the definition of functions $\mathcal{F}^{(i)}_{s,f}(\xi,\Omega)|_{p_{z}=p_{z,f}^{(s^{\prime})}}$, which define individual polarization components labeled by $i=0,+,-,\parallel$. Their explicit expressions are given in Eqs. (49) – (52) in Appendix E.
III Numerical results
@@
where we introduced the transverse momentum $k_{T}$, the azimuthal angle $\phi$, and the rapidity $y$, replacing the Cartesian components of the momentum, $k_{x}$, $k_{y}$, and $k_{z}$. By definition, $k_{y}=k_{T}\cos\phi$, $k_{z}=k_{T}\sin\phi$, and $y=\frac{1}{2}\ln\frac{k_{0}+k_{x}}{k_{0}-k_{x}}$. For simplicity, below we will be presenting the results for $y=0$ (i.e., $k_{x}=0$).
@@
III.1 Nonzero quark-number chemical potential $\mu$
@@
The corresponding numerical results are presented in Fig. 2, which confirm a substantial degree of circular polarization in the photon emission. This sign of ${\cal P}_{\rm circ}(k_{T},\phi,y)$ is determined primarily by a nonzero electrical charge density in the plasma. Indeed, by separating the partial contributions from the up and down quarks, we find that the up quarks emit more left-handed photons (${\cal P}^{\rm(u)}_{\rm circ}<0$), while the down quarks emit more right-handed photons (${\cal P}^{\rm(d)}_{\rm circ}>0$). Since the up quarks have twice the electric charge of the down quarks, their emission rate is higher overall. Consequently, the net circular polarization from QGP is negative, ${\cal P}_{\rm circ}<0$.
@@
III.2 Nonzero chiral chemical potential $\mu_{5}$
@@
Fig. 3 presents sample numerical results for the emission rates of left-handed and right-handed circularly polarized photons. The data is shown for a temperature of $T=200~{}\mbox{MeV}$ and for two different magnetic field strengths: $|eB|=m_{\pi}^{2}$ (left panel) and $|eB|=5m_{\pi}^{2}$ (right panel). Although we do not show the rates at $T=350~{}\mbox{MeV}$, we have verified that they exhibit a similar qualitative behavior, albeit with generally higher values.
@@
Figure 4: The emission asymmetry for the right-handed circularly polarized photons as a function of the azimuthal angular for two magnetic fields, $|eB|=m_{\pi}^{2}$ (left) and $|eB|=5m_{\pi}^{2}$ (right), and two temperatures, $T=200~{}\mbox{MeV}$ (top) and $T=350~{}\mbox{MeV}$ (bottom).
@@
III.3 Nonzero quark-number and chiral chemical potentials $\mu$ and $\mu_{5}$

### v2

# Circularly polarized photon emission from magnetized chiral plasmas

@@ I Introduction
High temperature and energy density are not the only extreme characteristics of the QGP produced by heavy-ion collisions. The system is, without exaggeration, rich in extremes. Since heavy ions carry electric charges, they induce extreme magnetic fields [1, 2, 3, 4, 5]. Furthermore, large values of angular momentum in noncentral collisions can produce high vorticity in the resulting QGP fluid [6, 7]. Many theoretical aspects of QGP associated with strong magnetic fields and high vorticity are routinely tested in experimental studies [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. (For reviews, see Refs. [22, 23, 24, 25, 26].)
One aspect we would like to discuss in more detail is the effect of chiral charge density on the electromagnetic probes of QGP. A nonzero chiral charge arises from an imbalance between left-handed and right-handed quarks in the plasma. At first glance, such a state may seem too exotic as it would break local parity symmetry. However, in a hot plasma, local bubbles of chiral plasma can be produced by topological configurations of gauge fields [27]. This idea is not new, and it is taken seriously enough to motivate the challenging search for its observable signatures via the dipole and quadrupole chiral magnetic effects (CME) in multiple experiments [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Unfortunately, the evidence supporting CME remains inconclusive.
@@
This paper is organized as follows. In Sec. II, we briefly present the formalism used for studying polarized photon emission from the plasma. The derivation of the polarization tensor for a plasma with nonzero quark number and chiral chemical potentials is outlined in Sec. III. Numerical results for the polarized photon emission are given in Sec. IV. In Sec. V, we discuss our main findings and summarize the results. Several technical derivations and auxiliary results are provided in the appendices at the end of the paper.
@@ II Formalism
At leading order in the coupling, photon emission from a hot, strongly magnetized QGP is primarily driven by quark and antiquark splitting processes: $q_{n}\to q_{n^{\prime}}+\gamma$ and $\bar{q}_{n}\to\bar{q}_{n^{\prime}}+\gamma$, as well as by quark-antiquark annihilation, $q_{n}+\bar{q}_{n^{\prime}}\to\gamma$, shown in Fig. 1. Here $q_{n}$ and $\bar{q}_{n^{\prime}}$ represent quark and antiquark states in the $n$th and $n^{\prime}$th Landau levels, respectively [31]. These processes contribute at the linear order in the fine structure constant $\alpha=1/137$, but they are forbidden by energy-momentum conservation in the absence of a magnetic field. When $B=0$, the dominant contribution instead comes from gluon-mediated two-to-two processes: $q+g\rightarrow q+\gamma$, $\bar{q}+g\rightarrow\bar{q}+\gamma$, and $q+\bar{q}\rightarrow g+\gamma$, where $g$ represents a gluon [32, 33, 34, 35, 36, 37, 38]. The rates of these processes are of the order of $\alpha\alpha_{s}$, where $\alpha_{s}$ is the strong coupling constant. Thus, when the one-to-two and two-to-one processes become significant in the presence of a strong magnetic field, the two-to-two processes yield only subleading corrections. Here we assume that the QGP is at a sufficiently high temperature, ensuring that the running value of $\alpha_{s}$ remains small. However, even in this case, the interplay between the two types of processes becomes highly nontrivial as the magnetic field strength decreases.
Figure 1: Three leading-order processes contributing to photon emission in a strongly magnetized QGP: (a) $q\to q+\gamma$, (b) $\bar{q}\to\bar{q}+\gamma$, (c) $q+\bar{q}\to\gamma$.
To investigate the impact of a nonzero chiral charge density on photon emission from a strongly magnetized plasma, we will utilize the general framework developed in Refs. [31, 39, 40], where unpolarized photon emission was studied. Specifically, we will utilize a relationship between the emission rate and the imaginary part of the retarded photon polarization tensor [41]
@@
Assuming the background magnetic field is sufficiently strong, the imaginary part of the polarization tensor $\Pi_{\mu\nu}(k)$ is determined by the simplest one-loop diagram at leading order in the coupling [39]. Since the branch cuts of the diagram are associated with intermediate two quark states going on the mass shell, $\mbox{Im}\left[\Pi_{\mu\nu}(k)\right]$ is proportional to the sum of squared amplitudes of the corresponding one-to-two and two-to-one processes, as depicted in Fig. 1. For the QGP, this leading-order approximation holds when the temperature is sufficiently high and the magnetic field is strong enough, specifically when $T\gg\Lambda_{\rm QCD}$ and $|eB|\gtrsim\alpha_{s}T^{2}$. A similar approximation applies to a QED plasma when $|eB|\gtrsim\alpha T^{2}$.
To separate total photon emission into polarized components, we use the following completeness relation:
$g^{\mu\nu}=\epsilon_{0}^{\mu}\epsilon_{0}^{\nu*}-\epsilon_{+}^{\mu}\epsilon_{+}^{\nu*}-\epsilon_{-}^{\mu}\epsilon_{-}^{\nu*}-\epsilon_{\parallel}^{\mu}\epsilon_{\parallel}^{\nu*},$ (2)
@@
In this study, we focus on the emission of real photons into the vacuum outside of QGP. Since the mean free path of photons is much larger than the system size in heavy-ion collisions, intermediate modes in the plasma can be ignored. Once produced, these modes are projected onto vacuum states without any decoherence from scattering. As long as the corresponding set of states is complete, the emission rate of vacuum photons remains effectively unchanged. Similar arguments were used for dilepton production in Ref. [43].
@@
Interestingly, the extra contribution ${\cal R}^{\rm(extra)}_{\rm diff}$ to the rate, coming from $\Pi^{\rm(extra)}(k)\equiv\Pi^{(0)}(k)+\Pi^{(\parallel)}(k)$, is negative. At first glance, this might seem problematic since observable rates, which are proportional to the positive-definite probability of emitting physical photons, are expected to always be positive. In vacuum, there are only two physical photon states with transverse polarizations. Therefore, the additional contributions come from unphysical degrees of freedom that are necessary to preserve gauge invariance in the theoretical framework used. However, as one can verify, the total rate ${\cal R}_{\rm diff}$ remains positive overall. This suggests that the extra contribution ${\cal R}^{\rm(extra)}_{\rm diff}$ must be distributed between the two physical circular polarizations, thereby modifying the rate definitions in Eqs. (5) and (6). For simplicity, we will continue using the unmodified definitions, which should be sufficient for approximately extracting parity-odd contributions to the rates for circularly polarized photons, albeit with a slight overestimation.
III Polarization tensor
@@
$\bar{G}_{f}(p_{0},p_{z};\mathbf{r}_{\perp})=i\frac{e^{-\mathbf{r}_{\perp}^{2}/(4\ell_{f}^{2})}}{2\pi\ell_{f}^{2}}\sum_{n=0}^{\infty}\frac{\tilde{D}_{n}^{f}(p_{0},p_{z};\mathbf{r}_{\perp})}{\mathcal{M}-2n|e_{f}B|},$ (7)
@@
Here, $L_{n}^{\alpha}(z)$ are the generalized Laguerre polynomials [44], $\mathcal{P}_{\pm}^{f}\equiv\frac{1}{2}\left(1\pm is_{\perp}\gamma^{1}\gamma^{2}\right)$ are the spin projectors, $\mathbf{r}_{\perp}=(x,y)$ is the position vector in the plane perpendicular to the field, and $\ell_{f}=1/\sqrt{|e_{f}B|}$ is the magnetic length. Here, we included nonzero quark number and chiral chemical potentials, denoted by $\mu$ and $\mu_{5}$, respectively. By definition, $s_{\perp}=\,\mbox{sign}(e_{f}B)$ and $L_{-1}^{\alpha}(z)\equiv 0$.
@@
where $\mathcal{P}_{5}^{(\pm)}=\frac{1}{2}(1\pm\gamma^{5})$ are the chirality projectors. Note that we included both the quark-number and chiral-charge chemical potentials, $\mu$ and $\mu_{5}$, respectively. Both play an important role in polarized photon emission. For the sake of simplicity, we assume that $\mu$ and $\mu_{5}$ are same for both flavors. (The generalization to the case of flavor-dependent chemical potentials is straightforward and will be briefly discussed later.)
@@
where we used the Matsubara finite-temperature formalism. By definition, the fermionic and bosonic Matsubara frequencies are $\omega_{k}=(2k+1)\pi T$ and $\Omega_{m}=2m\pi T$, respectively. Regarding other notations, $\alpha_{f}=q_{f}^{2}\alpha$ is the flavor-dependent coupling constant, and $N_{c}=3$ is the number of quark colors.
By substituting Green’s function (7) into the definition for $\Pi^{\mu\nu}(i\Omega_{m};\mathbf{k})$, calculating the Dirac traces, and finally integrating over $r_{\perp}$, we derive the following expression for the individual components of the polarization tensor
$\Pi^{(i)}(i\Omega_{m};\mathbf{k})=-\sum_{f=u,d}\frac{N_{c}\alpha_{f}T}{\pi\ell_{f}^{4}}\sum_{n,n^{\prime}=0}^{\infty}\sum_{k=-\infty}^{\infty}\sum_{s=\pm 1}\int\frac{dp_{z}}{2\pi}\frac{\mathcal{J}_{(i)}^{f}(-s,-s)}{[(i\omega_{k}+\mu+s\mu_{5})^{2}-E_{n,p_{z},f}^{2}][(i\omega_{k}+\mu-i\Omega_{m}+s\mu_{5})^{2}-E_{n^{\prime},p_{z}-k_{z},f}^{2}]},$ (11)
where $i$ labels the photon polarizations (i.e., $i=0,+,-,\parallel$) and $E_{n,p_{z},f}=\sqrt{p_{z}^{2}+2n|e_{f}B|}$ are Landau-level energies. The results for the Dirac traces are given in Appendix B and the explicit expressions for the functions $\mathcal{J}_{(i)}^{f}(-s,-s)$ are derived in Appendix C.
After performing the Matsubara summation, we derive the following final expressions for the polarization components:
@@
where $n_{F}(E)=1/\left[\exp(E/T)+1\right]$ is the Fermi-Dirac distribution function. The corresponding technical details as well as the explicit expressions for the functions $\mathcal{F}^{(i)}_{s,f}(\xi,i\Omega_{m})$ are given in Appendix D.
@@
where $\theta(x)$ is the Heaviside step function and $p_{z,f}^{(s^{\prime})}$ are the solutions of the energy conservation equation, given in Eq. (46) in Appendix D. In Eq. (13), we also introduced two transverse momentum thresholds,
$k_{\pm}^{f}=\left|\sqrt{2n|e_{f}B|}\pm\sqrt{2n^{\prime}|e_{f}B|}\right|,$ (14)
@@
As expected, the expression in Eq. (13) closely resembles the result in Ref. [39]. It should be noted, however, that $g(n,n^{\prime})$ depends on both fermion-number and chiral chemical potentials. The most important difference appears in the definition of functions $\mathcal{F}^{(i)}_{s,f}(\xi,\Omega)|_{p_{z}=p_{z,f}^{(s^{\prime})}}$, which define individual polarization components labeled by $i=0,+,-,\parallel$. Their explicit expressions are given in Eqs. (50) – (53) in Appendix D.
IV Numerical results
@@
where we introduced the transverse momentum $k_{T}$ (lying in the plane perpendicular to the ion beams), the azimuthal angle $\phi$, and the rapidity $y$, replacing the Cartesian components of the momentum, $k_{x}$, $k_{y}$, and $k_{z}$. By definition, $k_{y}=k_{T}\cos\phi$, $k_{z}=k_{T}\sin\phi$, and the rapidity $y=\frac{1}{2}\ln\frac{k_{0}+k_{x}}{k_{0}-k_{x}}$. For simplicity, below we will be presenting the results for zero rapidity (i.e., $k_{x}=0$). The illustration of the coordinate system used is shown in Fig. 2.
Figure 2: The illustration of the coordinate system used, where the magnetic field $\mathbf{B}$ is perpendicular to the reaction plane. The reaction plane is parameterized by $\mathbf{r}_{\perp}=(x,y)$. The photon momentum $\mathbf{k}_{T}=k_{T}(0,\cos\phi,\sin\phi)$ lies in the transverse $yz$-plane, which is perpendicular to the beam.
@@
IV.1 Nonzero quark-number chemical potential $\mu$
@@
As we see, the differential rates exhibit nonsmooth behavior as a function of the angular coordinate $\phi$. They are characterized by numerous spikes resulting from the Landau-level quantization of quark states. This prominent feature was discussed in detail in Ref. [39]. Strictly speaking, the emergence of threshold spikes is an artifact of the approximation that neglects the self-energy of quarks. All threshold singularities would smooth out if the quark damping rates, $\Gamma_{n}(p_{z})$, were considered. Notably, these damping rates were recently calculated in the Landau-level representation in Ref. [45]. Assuming small $\Gamma_{n}(p_{z})$, the corresponding modifications occur only within narrow energy windows near the thresholds. Overall, the qualitative effect of nonzero $\Gamma_{n}(p_{z})$ is expected to result in localized smoothing while preserving the global features of the rates [39]. Thus, for the purposes of the current study, the inclusion of quark damping rates is not critical.
@@
The corresponding numerical results are presented in Fig. 4, which confirm a substantial degree of circular polarization in the photon emission. The sign of ${\cal P}_{\rm circ}(k_{T},\phi,y)$ is determined primarily by a nonzero electrical charge density in the plasma. Indeed, by separating the partial contributions from the up and down quarks, we find that the up quarks emit more left-handed photons (${\cal P}^{\rm(u)}_{\rm circ}<0$), while the down quarks emit more right-handed photons (${\cal P}^{\rm(d)}_{\rm circ}>0$). Since the up quarks have twice the electric charge of the down quarks, their emission rate is higher overall. Consequently, the net circular polarization from QGP is negative, ${\cal P}_{\rm circ}<0$.
@@
IV.2 Nonzero chiral chemical potential $\mu_{5}$
@@
Fig. 5 presents sample numerical results for the emission rates of left-handed and right-handed circularly polarized photons. The data is shown for a temperature of $T=200~{}\mbox{MeV}$ and for two different magnetic field strengths: $|eB|=m_{\pi}^{2}$ (left panel) and $|eB|=5m_{\pi}^{2}$ (right panel). Although we do not show the rates at $T=350~{}\mbox{MeV}$, we have verified that they exhibit similar qualitative behavior, albeit with generally higher values.
@@
Figure 6: The emission asymmetry for the right-handed circularly polarized photons as a function of the azimuthal angle for two magnetic fields, $|eB|=m_{\pi}^{2}$ (left) and $|eB|=5m_{\pi}^{2}$ (right), and two temperatures, $T=200~{}\mbox{MeV}$ (top) and $T=350~{}\mbox{MeV}$ (bottom).
@@
IV.3 Nonzero quark-number and chiral chemical potentials $\mu$ and $\mu_{5}$